OpenStudy (anonymous):

Anyone good at anything I could really use help !! Trying to get into uni and have to sit a math test havent been in school properly for two years since I was in year ten. Please, I'm begging youu !!

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):

\[{20-2[3+(12/4)]}\]

OpenStudy (anonymous):

So lost ..

OpenStudy (unklerhaukus):

ok so the first step , is to simplify the fraction

OpenStudy (anonymous):

which can simplify to 6/1 ?

OpenStudy (unklerhaukus):

hmm, not quite

OpenStudy (anonymous):

9/3 ?

OpenStudy (anonymous):

Wait, wait, wait. Let me introduce you to my friend PEMDAS. He is an expert in mathematics.

OpenStudy (unklerhaukus):

\[12=4\times3\] \[\frac{12}4=\frac{4\times3}4=\frac{\cancel{4\times}3}{\cancel{4}}=\frac 31=3\]

OpenStudy (anonymous):

I'm sorry but I dont understand that. The only math I have done in the last two years is counting money for work. I get the first line. 12 goes into four three times but the rest you have lost me.

OpenStudy (unklerhaukus):